Business production: A business has three machines that manufacture containers. Together they make 100 containers per day, whereas the two fastest machines can make 80 containers a day. The fastest machine makes 34 more containers per day than the slowest machine.
A. letting x, y, and z, be the number of containers that the machines make from fastest to slowest, write a system of three equations whose solution gives the number of containers that each machine can make.
B. Solve the system

Answer by jorel1380(3719)

x+y+z=100
x+y=80
x-z=34
-------------
x-34=z
(x+y)+z=100
80+x-34=100
x=54
y=80-54=26
z=x-34=20
Machines x,y,and z make 54,26, and 20 containers a day, respectively..